OpenStudy (anonymous):

A spherical balloon is being inflated. The radius of the balloon is increasing at the rate of 6 cm/s. f(t) = 6t Find a function g that models the volume as a function of the radius r. g(r) = ?

OpenStudy (turingtest):

so r=f(t) then...

OpenStudy (turingtest):

what is the formula for the volume of a sphere?

OpenStudy (anonymous):

Are you solving this by using derivatives?

OpenStudy (anonymous):

Dr/Dt= 6cm/s V=4/3 π r3 Dv/Dt=4πr^2(Dr/Dt) Plug in your variables: Dv/Dt=4πr(6) So the function would be: Dv/Dt=24πr